# T3-S2S: Training-free Triplet Tuning for Sketch to Scene Generation

### v1

@@ 4.1 Overview
(2) Characteristics Prominence: This module selects instance-related tokens and their sketches by identifying the TopK values for each channel in the value matrices, creating an instance-specific mask. The mask is then used to scale up the feature map for the corresponding channel. This approach enhances the distinction of instances within the multi-channel feature space without additional parameters, ensuring that instances’ characteristics are more prominently emphasized.
(3) Dense Tuning: While prompt balance increases the strength of the embedding matrices related to instances, enhancing their competitiveness in the attention map, the overall strength of the attention map remains suboptimal. Meanwhile, given that more contour information resides in the ControlNet branch, we employ dense modulation directly within this branch to augment the attention map for better modulation. Specific implementation refers to Dense Diffusion [4].
Building on these three modules, a unified training-free triplet tuning strategy is implemented throughout the entire cross-attention mechanism. This ensures that the final generation effectively responds to both text and sketch inputs, thereby enhancing the stability and diversity of the generated outputs. In the subsequent section, we will provide a detailed explanation of two newly designed modules and their underlying rationales.
Figure 5: Overview of the proposed training-free triplet tuning strategy in the frozen pre-trained latent diffusion model. (a) The orange parts indicate the proposed module plugged into the ControlNet and U-Net framework. (b) The left part shows the energy tuning of prompt balance. (c) The bottom part indicates the training-free tuning of characteristics prominence.
@@ 4.2 Prompt Balance
As discussed in Section 3.2, the imbalance of prompt energy related to instances influences representations of key and value matrices, and causes instances missing in the generated image. Therefore, by enhancing the balance of keywords’ energy and scaling up the values, we can improve the instance accuracy and details in the generated images. To do this, we propose a plug-in strategy for the text embeddings shown in Figure 5 (b), named prompt balance.
Specifically, we use an NLP network (e.g., the SpaCy library) to identify instance keywords from the global text prompts $\mathbf{c}_{g}=\{c^{i}\}^{l}_{i=0}$, resulting in reorganized instance keyword prompts $\{c^{i}\}^{\mathbf{q}}$, where $\mathbf{q}$ is the indices vector of keywords in $\mathbf{c}_{g}$. Then, we encode both the global text prompts and each instance keyword prompts separately into text embeddings $\mathbf{S}_{g}=\{\mathbf{s}^{i}_{g}\}\in\mathbb{R}^{n\times d}$ and $\{\mathbf{s}^{i}_{w}\in\mathbb{R}^{1\times d}\}^{\mathbf{q}}$ by a text encoding network. Next, we replace the embedding of keywords in $\mathbf{S}_{g}$ with the single-word embedding of $\mathbf{S}_{w}$ to form a new combined embedding $\mathbf{S}_{r}\in\mathbb{R}^{n\times d}$, defined as:
@@
Generally, the special “end of text” token (located $i_{\text{end}}$) always has the maximum energy as shown in Section 3.2, which could be the upper bound for us to scale up the embeddings of the keywords in $\mathbf{S}_{r}$ that all keywords have balanced energy relative to the “end of text” token embedding, mathematically represented as:
@@
Finally, the balanced text embeddings, denoted as $\mathbf{S}_{b}$, will benefit the values of instance-based tokens in the key and value matrices, as well as the attention map. Subsequently, it will enhance the competitiveness of instance tokens among all tokens while maintaining consistency across instance tokens, providing a concise summary that highlights the importance of each instance.
Figure 6: Qualitative comparison with baseline methods. (a) T3-S2S performs well for smaller instances like “houses” and “path” , and unusual “mountain”. (b) T3-S2S performs well with a large number of small instances “trees”. (c) T3-S2S decouples the overlap of instances. Note that the original Dense Diffusion [4] based on SD V1.5 [1], has limited prompt response capabilities. For a fair comparison, we apply it to the SDXL model.
4.3 Characteristics Prominence
While utilizing balanced text embeddings helps balance the competition among instances, diffusion models still face challenges with entity coupling in cross-attention layers, lacking a mechanism to address the interaction between the attention map and value matrices. As discussed in Section 3.3, increasing the TopK values in each channel of the value matrices reduces value homogeneity, but requires a trade-off between instance completeness and noise clarity. In this subsection, we introduce a characteristics prominence technique, applied after feature map computation in the original cross-attention layers, without introducing any additional trainable parameters.
Specifically, instead of directly enhancing the TopK values along the $n$ dimension in the value matrix $\mathbf{V}_{m}\in\mathbb{R}^{n\times d_{m}}$, we apply enhancement based on the indices of the TopK values on the feature map $\mathbf{F}_{m}\in\mathbb{R}^{{b_{m}}\times d_{m}}$ (before the residual adding). For each channel in the value matrix $\mathbf{V}_{m}$, we find the indices of the TopK values across all valid tokens (between “start of text ” and “end of text” tokens):
@@ 4.3 Characteristics Prominence
$\hat{\mathbf{F}}_{m}=\mathbf{F}_{m}+\beta\cdot\mathbf{H}_{m}\odot\mathbf{F}_{m},$
where $\odot$ denotes element-wise multiplication. This enhancement emphasizes instance tokens within the multi-channel feature space, aiding in distinguishing each instance more effectively. The characteristics prominence technique strengthens the attention mechanism by ensuring that each instance is highlighted, even when its sketch is small. By amplifying relevant regions in the feature map, the model improves instance differentiation, making it better suited for multi-instance scene generation.
@@ 5.1 Implementation Details
Baselines. We leverage the sketch-processing capabilities of the SDXL-base model [2] and the ControlNet model [3, 14], serving as our foundational models. Additionally, we extend our comparison to include two sketch-oriented approaches: the training-based T2I Adapter [30] and the training-free Dense Diffusion [4], both integrated with the SDXL-base model.
Setup. In the triplet tuning scheme, the prompt balance module is integrated into the text encoding process, while the characteristics prominence modules are incorporated across all cross-attention layers. Additionally, the dense tuning module is specifically added to the “down_blocks 2” layers and the “mid_blocks 0” layers within the ControlNet branch. The TopK value $K$ is set to 2, and $\beta$ is kept at 1. During inference, we use the default Euler Discrete Scheduler [39] with 32 steps and a guidance scale of 9 at a resolution of $1024\times 1024$. All experiments are conducted on a single Nvidia Tesla V100 GPU.
Metrics. Given that our current approach involves sketch-based multi-instance scene generation, existing benchmarks should be adjusted for our evaluation, such as adding sketch inputs for T2I-CompBench [40]. Therefore, we design 20 complex sketch scenes, each with more than four sub-prompts, encompassing various terrains (plains, mountains, deserts, tundra, cities) and diverse instances (rivers, bridges, stones, castles). We utilize CLIP-Score [41] for the global prompt and image, and evaluate the CLIP-Score for each background prompt and instance prompts by cropping the corresponding regions. Additionally, we conduct a user study to assess different variants of our approach, using a 1-5 rating scale to evaluate image quality, placement, and prompt-image consistency. Details can be found in Appendices C and  E.
@@ 5.2 Main Results
Qualitative Evaluation. Building upon the scene design, we develop three representative and complex multi-instance scene scenarios, each incorporating a diverse array of elements to foster varied interactions. We evaluate several approaches, with visual comparisons displayed in Figure 6. Due to the specialized nature of this task, most existing solutions are inadequate, often overlooking small objects and failing to manage instance overlap effectively. When combined with the triplet tuning strategy, our T3-S2S method improves the generation performance of existing SDXL models. For example, Figure 6 (a) showcases the enhanced detail in smaller instances such as “houses” and “path”, and even less common elements like “mountains”. Similarly, Figure 6 (b) illustrates the effective generation of numerous small instances, like “trees”. Furthermore, our approach excels in scenarios with complex instance contour interactions, as depicted in Figure 6 (c), accurately capturing and displaying all details. By leveraging the triplet tuning strategy and an advanced cross-attention mechanism, our approach consistently generates detailed, multi-instance scenes that closely adhere to the original sketches and prompts, ensuring both stability and diversity in generations. Additional game scenes and diverse scenes are provided in Appendices C and  D.
Table 1: Comparison of CLIP-Score across several variants, evaluated on whole images, masked instance regions, and masked background regions. Includes user study ratings on a scale of 1-5.
@@
ControlNet 0.3440 0.2423 0.2539 2.34
PB 0.3447 0.2479 0.2562 2.64
DT 0.3433 0.2463 0.2549 3.29
CP 0.3467 0.2490 0.2563 3.45
PB+CP 0.3465 0.2548 0.2573 3.61
T3-S2S 0.3497 0.2563 0.2588 3.88
Quantitative Evaluation. We compare CLIP-Scores for global image, instances, and background across different variants and the base ControlNet. A user study is also conducted with a 1-5 rating scale. As shown in Table 1, our approach demonstrates superior performance on the 20 complex multi-instance scenes, with improved fidelity and precision in aligning with text prompts and sketch layouts. The PB module shows modest improvement, while the CP and DT modules provide significant and comparable enhancements. Combining these components allows our T3-S2S approach to achieve a well-balanced outcome.
Figure 7: Visual comparison of different inserted modules. DT: Dense Tuning; PB: Prompt Balance; CP: Characteristics Prominence.
@@ 5.3 Ablation Study
Module Comparison. In this validation, we conduct an ablation study to assess the individual and combined impacts of different modules, with the findings detailed in Figure 7. While each module contributes to improving generation quality, no single module fully resolves all challenges: (1) Dense tuning effectively restricts instance overlap within sketch areas, such as “bridges”, by optimizing instance overlap. (2) Prompt balance enhances the visibility of smaller objects like “houses”, although it may inadvertently introduce noise associated with these houses. (3) Characteristic prominence sharpens the distinct features of instances, enhancing clarity and reducing irrelevant noise. (4) A combined application of prompt balance and characteristic prominence effectively addresses most issues, approaching success. When these three modules are integrated to form our triplet tuning approach, they enhance the alignment between generated scene images and their corresponding sketches and prompts, leading to more consistent and accurate representations. To further verify the functions of modules, we transfer the PB module to the Attend-and-Excite [42] in Appendix F, and the T3-S2S to the T2I adapter [30] in Appendix G.
Figure 8: Visual comparison of two hyper-parameters $K$ and $\beta$ suggests that setting $K=2$ and $\beta=1$ is a favorable choice.
Hyper-parameter Comparison. To validate our hypothesis, we examine the impacts of varying $K$ and $\beta$ values on generations, with visual results presented in Figure 8. In a set of experiments where $\beta$ is fixed at 1, we find that increasing $K$ initially improves generation quality but eventually leads to more noise. Conversely, when $K$ is held steady at 2, adjusting $\beta$ above 1 consistently produces favorable outcomes, maintaining stable generation quality across higher $\beta$ values. Based on these observations, we determine that the optimal settings for our model are $K=2$ and $\beta=1$.We also conduct an analysis of TOP $K$ distribution in Appendix B.

### v2

@@ 4 Diffusion for Controllable Concept Art Generation
(2) Characteristics Priority: This module selects instance-related tokens and their sketches by identifying the TopK values for each channel in the value matrices, creating an instance-specific mask. The mask is then used to scale up the feature map for the corresponding channel. This approach enhances the distinction of instances within the multi-channel feature space without additional parameters, ensuring that instances’ characteristics are more prominently emphasized.
(3) Dense Tuning: While prompt balance increases the strength of the embedding matrices related to instances, enhancing their competitiveness in the attention map, the overall strength of the attention map remains suboptimal. Meanwhile, given that more contour information resides in the ControlNet branch, we employ dense modulation directly within this branch to augment the attention map for better modulation. Since attention manipulation has been extensively explored in prior works (Hertz et al., 2022; Xie et al., 2023; Kim et al., 2023; Chen et al., 2024), we only adopt a modified dense tuning scheme from Kim et al. (2023), detailed in Section 3.
In the subsequent section, we will provide a detailed explanation of two newly designed modules and their underlying rationales.
Figure 6: Overview of the proposed training-free triplet tuning strategy in the frozen pre-trained latent diffusion model. (a) The orange parts indicate the proposed module plugged into the ControlNet and U-Net framework. (b) The left part shows the energy tuning of the prompt balance. (c) The bottom part indicates the training-free tuning of the characteristics priority. T3-S2S  ensures the final generation responds effectively to both text and sketch inputs.
@@ 4.1 Prompt Balance
As discussed previously, the imbalance of prompt energy influences competition among key and value matrices related to instances, increasing the risk of missing instances. To mitigate this, we propose a plug-in prompt balance strategy for the text embeddings that enhances energy uniformity across keywords and scales up values, as displayed in Figure 6 (b). Specifically, we use an NLP network (e.g., the SpaCy library) to identify instance keywords from the global text prompts $\mathbf{c}_{g}=\{c^{i}\}^{l}_{i=0}$, resulting in reorganized instance keyword prompts $\{c^{i}\}^{\mathbf{q}}$, where $\mathbf{q}$ is the indices vector of keywords in $\mathbf{c}_{g}$. Then, we encode both the global text prompts and each instance keyword prompts separately into text embeddings $\mathbf{S}_{g}=\{\mathbf{s}^{i}_{g}\}\in\mathbb{R}^{n\times d}$ and $\{\mathbf{s}^{i}_{w}\in\mathbb{R}^{1\times d}\}^{\mathbf{q}}$ by a text encoding network. Next, we replace the embedding of keywords in $\mathbf{S}_{g}$ with the single-word embedding of $\mathbf{S}_{w}$ to form a new combined embedding $\mathbf{S}_{r}\in\mathbb{R}^{n\times d}$:
@@
Generally, the special “end of text" token (located $i_{\text{end}}$) always has the maximum energy as shown in Fig. 3, which could be the upper bound for us to scale up the embeddings of the keywords in $\mathbf{S}_{r}$ that all keywords have balanced energy relative to the “end of text" token embedding, mathematically represented as:
@@
Finally, the balanced text embeddings, denoted as $\mathbf{S}_{b}$, enhance the instance-based token values in the key, value matrices, and attention map, improving their competitiveness and consistency for more concise and effective instance representation.
4.2 Characteristics Priority
While balanced text embeddings help equalize instance competition, value matrices still face low prominence. As discussed in Section 3, enhancing TopK values per channel improves prominence. To balance instance completeness and noise clarity, we propose a CP technique that applies localized TopK enhancement to specific sketch regions for features. Specifically, instead of directly enhancing the TopK values along the $n$ dimension in the value matrix $\mathbf{V}_{m}\in\mathbb{R}^{n\times d_{m}}$, we apply enhancement based on the indices of the TopK values on the feature map $\mathbf{F}_{m}\in\mathbb{R}^{{b_{m}}\times d_{m}}$ (before the residual adding). For each channel in the value matrix $\mathbf{V}_{m}$, we find the indices of the TopK values across all valid tokens (between “start" and “end" tokens):
@@ 4.2 Characteristics Priority
$\mathbf{\hat{F}}_{m}=\mathbf{F}_{m}+\beta\cdot\mathbf{H}_{m}\odot\mathbf{F}_{m},$
where $\odot$ denotes element-wise multiplication. This enhancement emphasizes instance tokens in the multi-channel feature space, improving instance distinction. The CP technique strengthens the attention mechanism by amplifying instance-relevant regions in the feature map, even with small sketches.
4.3 Structural Concept Representations
Building on the previous modules, the T3-S2S  can produce detailed, multi-instance scenes that align closely with input prompts and sketches. To further enhance scene structure, especially in maintaining terrain consistency while separating foreground objects, we propose a twin structural concept representations framework. As depicted in Figure 7, it adopts a dual-branch prompt synthesis with mask-guided feature-sharing mechanism: one for the complete isometric image with full sketches and another for the empty terrain with non-foreground sketches.
Figure 7: Twin-structure framework with dual-branch prompt synthesis, where a mask-guided feature-sharing mechanism handles full-sketch isometric and non-foreground terrain images.
The feature maps for the complete isometric image and the empty terrain are denoted as $\mathbf{F}_{\text{iso}}(t)$ and $\mathbf{F}_{\text{terr}}(t)$, respectively, at inference step $t$. The terrain branch updates with selective and progressive blending, defined as:
$\mathbf{\hat{F}}_{\text{ter}}=\mathbf{F}_{\text{ter}}+\mathbf{M}\odot\gamma(t)\cdot\big(\mathbf{F}_{\text{iso}}-\mathbf{F}_{\text{ter}}\big),\quad\gamma(t)=\gamma_{0}\cdot(\frac{t}{T})^{\alpha},$
where $\odot$ denotes element-wise multiplication, $\gamma_{0}$ the initial scaling factor. $\mathbf{M}$ is a binary mask derived from the sketch sub-prompts, with $1$ indicating background regions and 0 indicating foreground regions (e.g., houses, bridges). The dynamic scaling $\gamma(t)$ ensures that the terrain branch inherits global features from the isometric branch at the early stages, while allowing independent refinement later, with the mask $\mathbf{M}$ preserving foreground integrity. Together, mask-guided sharing and dynamic scaling enable harmonious terrain generation and effective foreground-background separation.
@@ 5.1 Implementation Details
Baselines. Our baseline leverages the sketch-processing capabilities of the ControlNet model (Zhang et al., 2023b; Xinsir, 2023) with the SDXL-base model (Podell et al., 2023), compared with two SDXL-base sketch-oriented approaches: the training-based T2I Adapter (Mou et al., 2023) and the training-free Dense Diffusion (Kim et al., 2023). We also further validate PB with Attend-and-Excite (Chefer et al., 2023) (Appendix G), and T3-S2S with T2I adapter (Appendix H).
Setup. In the triplet tuning scheme, the prompt balance module is integrated into the text encoding process, while the characteristics priority module are incorporated across all cross-attention layers. The dense tuning module is specifically added to the “down_blocks 2" layers and the “mid_blocks 0" layers within the ControlNet branch. We set $K=2$, $\alpha=5$, $\gamma_{0}=1$ and $\beta=1$. During inference, we use the default Euler Discrete Scheduler (Karras et al., 2022) with 32 steps and a guidance scale of 9 at a resolution of $1024\times 1024$. All experiments are conducted on a single Nvidia Tesla V100 GPU.
Metrics. Given that our current approach involves sketch-based multi-instance scene generation, existing benchmarks should be adjusted for our evaluation, such as adding sketch inputs for T2I-CompBench (Huang et al., 2023). Therefore, we design 50 complex sketch scenes, each with more than four sub-prompts, encompassing various terrains (plains, mountains, deserts, tundra, cities) and diverse instances (rivers, bridges, stones, castles). We utilize CLIP-Score (Hessel et al., 2021) for the global prompt and image, and evaluate the CLIP-Score for each background prompt and instance prompts by cropping the corresponding regions. Additionally, we conduct a user study to assess different variants of our approach, using a 1-5 rating scale to evaluate image quality, placement, and prompt-image consistency. Details can be found in Appendix D&F.
Figure 8: Qualitative comparison with baseline methods. (a) T3-S2S performs well for smaller instances like “houses" and “path" , and unusual “mountain". (b) T3-S2S performs well with a large number of small instances “trees". Note that the original Dense Diffusion (Kim et al., 2023) based on SD V1.5 (Rombach et al., 2022), has limited prompt response capabilities. For a fair comparison, we apply it to the SDXL model.
Figure 9: Qualitative comparison of different inserted modules based on the ControlNet and SDXL model. The quantitative results are present in Table 2.
@@ 5.2 Main Results
Qualitative Evaluation. Building upon the scene design, we show two representative and complex multi-instance scene scenarios, each incorporating a diverse array of elements to foster varied interactions. We evaluate several approaches, with visual comparisons displayed in Figure 9. Due to the specialized nature of this task, most existing solutions will overlook objects aligned with the prompts. When combined with the triplet tuning strategy, our T3-S2S method improves the generation performance of existing SDXL models. For example, Figure 9 (a) showcases the enhanced detail in smaller instances such as “houses" and “path", and even less common elements like “mountains". Similarly, Figure 9 (b) illustrates the effective generation of numerous small instances, like “trees". By leveraging the triplet tuning strategy within the cross-attention mechanism, our approach consistently generates detailed, multi-instance scenes that closely adhere to the original sketches and texts. Additional game and common scenes are provided in Fig. 15 and Appendix D&E.
Table 1: Evaluation results on T2I-CompBench. # indicates the method evaluated in the new benchmark, others from the T2i-CompBench paper.
Color 2D Spatial Numeracy
@@
SDXL 0.5879 0.2133 0.4991
ControlNet# 0.5925 0.2686 0.4639
ControlNet+DT# 0.6154 0.3419 0.5143
T3-S2S# 0.6872 0.4256 0.6257
Stable v3 0.8132 0.3200 0.6174
FLUX.1 0.7407 0.2863 0.6185
Table 2: Comparison of CLIP-Score across several variants based on SDXL model, evaluated on whole images, masked instance regions, and masked background regions. Includes user study ratings on a scale of 1–5.
[…]
DenseDiff 0.3438 0.2459 0.2543 2.25
T2I-Adaptor 0.3422 0.2405 0.2523 2.17
ControlNet 0.3435 0.2427 0.2531 2.38
+PB 0.3439 0.2472 0.2559 2.62
+DT 0.3430 0.2469 0.2547 3.18
+CP 0.3457 0.2503 0.2565 3.40
+PB+CP 0.3466 0.2552 0.2571 3.55
T3-S2S 0.3491 0.2566 0.2581 3.79
Quantitative T2I-CompBench. We evaluate common scene generation on the large-scale T2I-CompBench (Huang et al., 2023) using Color (B-VQA), 2D Spatial (UniDet), and Numeracy (UniDet) metrics (Standard: 300 prompts/metric, 10 images/prompt). As our method requires sketch inputs, we use GPT-4V to generate square sketches (similar to bounding boxes) for all instances, with manually adjusted sizes, presented in Table 2. Our method improves about 9.5% in the Color metric and 15.7% in the 2D Spatial metric over the ControlNet baselines. Especially for the Numeracy metric, we ensure that the 100 examples contain 1-3 small instance sketches with a minimal resolution of (ensuring 1 pixel at the minimal feature level). Compared to the SDXL baseline, ControlNet performs 3.5% drop on the Numeracy metric featuring small instances, while DenseDiffusion improves 1.4%. Our approach shows noticeable 16.2% gains over ControlNet for the generation of small instances.
Quantitative Customized Evaluation. We compare CLIP-Scores for global image, instances, and background across different variants and the base ControlNet. A user study is also conducted with a 1-5 rating scale. As shown in Table 2, our approach demonstrates better performance on the 50 complex multi-instance scenes, with improved fidelity and precision in aligning with text prompts and sketch layouts. The PB module shows modest improvement, while the CP and DT modules provide comparable enhancements. Combining these components allows our T3-S2S approach to achieve a well-balanced outcome.
@@ 5.3 Ablation Study
Module Comparison. We perform an ablation study (Fig. 9) to evaluate the individual and combined effects of three modules, as well as the quantitative results in Table 2: (1) DT constrains instance regions within sketches; (2) PB improves small object visibility (e.g., “houses”) but may introduce noise; (3) CP sharpens features and suppresses noise. PB+CP addresses most issues, and combining all three (T3-S2S) yields the best sketch-prompt alignment between generated scene images and their corresponding sketches and texts, aligning with the quantitative results present in Table 2.
(a) The first and second extreme points.
(b) The third and fourth extreme points.
Figure 10: Histogram of the distribution of extrema in Value matrices. At the first and second extrema, the probability of later instance tokens is lower than that of earlier tokens, so increasing their values helps enhance their representations. However, at the third and fourth extremes, the probabilities tend to converge, implying that only a few tokens are key to defining features.
Figure 11: Visual comparison of two hyper-parameters $K$ and $\beta$ suggests that setting $K=2$ and $\beta=1$ is a favorable choice. With $\beta=1$, increasing $K$ initially improves results but eventually adds noise. Fixing $K=2$, higher $\beta$ values yield stable and favorable outcomes. Another comparison is presented in Figure 18.
5.4 Top $K$ Anaylsis
In the module of characteristic prominence, two hyperparameters, $K$ and $\beta$, require meticulous tuning. $K$ determines the indices of extreme values within the value matrices. For our analysis, we save these indices and construct a histogram, as depicted in Fig. 10. We observe that the probability of later instance tokens achieving the maximum and second maximum values is comparatively lower than that of earlier tokens. Thus, increasing the value of later instance tokens will be beneficial for their representations. However, at the third and fourth extremes, the probabilities tend to converge, indicating that not every token is essential for defining key characteristics. Increasing values for later instances at this point would introduce additional noise. Therefore, setting $K=2$ is advisable based on the observed trends. For $\beta$, which enhances the characteristics of instances within the feature matrix, an initial increase is beneficial. Nonetheless, there is a critical threshold beyond which increases in $\beta$ begin to disrupt the distribution within the value matrices.
Hyper-parameter Comparison. To validate our hypothesis, we study TopKK distribution in Fig. 10. In Fig. 10, later tokens have lower chances of reaching top values at the first and second extrema, so increasing their values helps enhance their representations. But at the third and fourth extrema, probabilities even out, implying only a few tokens are key to defining features. Based on this observation, we further examine the effects of varying $K$ and $\beta$ on generation quality in Fig. 11. When fixing $\beta=1$, increasing $K$ initially improves visual quality by strengthening under-represented tokens, but further increasing $K$ introduces noticeable noise, as less relevant tokens are also amplified. Conversely, when fixing $K=2$, larger $\beta$ values lead to more stable and consistently favorable results, indicating that moderate amplification of a small number of key tokens is sufficient to achieve effective feature modulation without disrupting overall coherence. These results suggest a trade-off between expressiveness and stability, where overly large $K$ increases noise, while appropriately scaled $\beta$ preserves structural consistency. Consequently, we identify $K=2$ and $\beta=1$ as a robust and effective configuration across our experiments.